# Towards Principled Evaluations of Sparse Autoencoders for Interpretability and Control

### v2

@@ Abstract
Disentangling model activations into meaningful features is a central problem in interpretability. However, the lack of ground-truth for these features in realistic scenarios makes the validation of recent approaches, such as sparse dictionary learning, elusive. To overcome this, we propose a framework to evaluate feature dictionaries in the context of specific tasks, by comparing them against supervised feature dictionaries. First, we demonstrate that supervised dictionaries achieve excellent approximation, control and interpretability of model computations on the task. Second, we use the supervised dictionaries to develop and contextualize evaluations of unsupervised dictionaries along the same three axes.
We apply this framework to the indirect object identification task (IOI) using GPT-2 Small, with sparse autoencoders (SAEs) trained on either the IOI or OpenWebText datasets. We find that these SAEs capture interpretable features for the IOI task, but they are not as successful as supervised features in controlling the model. Finally, we observe two qualitative phenomena in SAE training: feature occlusion (where a causally relevant concept is robustly overshadowed by even slightly higher-magnitude ones in the learned features), and feature over-splitting (where binary features split into many smaller features without clear interpretation). We hope that our framework will be a useful step towards more objective and grounded evaluations of sparse dictionary learning methods.
@@
Figure 1: Overview of our evaluation pipeline. We start by choosing a specific model capability, and disentangling model activations into capability-relevant features using supervision. Then, we evaluate a given feature dictionary w.r.t. this capability, using the supervised features as a benchmark. We test the degree to which (1) the feature dictionary’s reconstructions of the activations are necessary/sufficient for the capability, (2) the features can be used to edit capability-relevant information in internal model representations (agnostic of feature interpretations), and (3) the features can be interpreted w.r.t. the capability in a way consistent with their causal role.
@@ 1 Introduction
While large language models (LLMs) have demonstrated impressive (Vaswani et al., 2017; Devlin et al., 2019; Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) results, the inner mechanisms behind their successes and failures largely remain a mystery (Olah, 2023). A central problem in this area is how to disentangle internal model representations into meaningful concepts, or features. If successful at scale, this research would deliver significant scientific and practical value, enabling enhanced model robustness, controllability, interpretability, and debugging (Gandelsman et al., 2023; Nanda et al., 2023; Marks et al., 2024).
@@
A leading hypothesis for how LLMs represent and use features is the linear representation hypothesis (Mikolov et al., 2013b; Grand et al., 2018; Li et al., 2021; Abdou et al., 2021; Nanda et al., 2023), a strong version of which posits that individual activations of a model can be decomposed as sparse linear combinations of features that come from a large, shared feature dictionary. Recently, a series of works have proposed applying the (unsupervised) sparse autoencoder (SAE) framework to find such dictionaries (Olshausen & Field, 1997; Faruqui et al., 2015; Goh, 2016; Arora et al., 2018; Yun et al., 2021; Cunningham et al., 2023; Bricken et al., 2023).
@@
While there are promising initial results (Bricken et al., 2023), this research area faces a key obstacle: we cannot directly evaluate the usefulness of features learned by an SAE, as we do not know the hypothetical ‘true’ features to begin with; indeed, finding them is why we use SAEs in the first place. This has two consequences: (1) the training objective of an SAE – balancing $\ell_{2}$ reconstruction with an $\ell_{1}$ penalty on feature coefficients – is only a substitute for the true goal of finding meaningful features, and (2) the metrics used to evaluate SAEs are indirect and rely on proxies for the features, toy models, or non-trivial assumptions on SAE learning (Elhage et al., 2022b; Bricken et al., 2023; Sharkey et al., 2023). This holds back progress in the field, as there is no simple, objective and direct way to compare trained SAEs. In this paper, we take steps towards addressing this challenge. Our contributions are outlined as follows:
@@
  • We propose a principled method to compute sparse feature dictionaries to decompose language model activations on realistic tasks, using supervision from labels of concepts describing the prompts.
@@
  • We apply this method to the IOI task, and show these dictionaries exhibit three desirable properties in the task’s context: (1) sufficiency/necessity of activation reconstructions, (2) sparse controllability of model behavior via feature editing, and (3) interpretability of the features consistent with their causal role000 As a by-product, our supervised feature dictionaries also demonstrate that activations for the task can be usefully disentangled in a way that adheres to the linear representation hypothesis and exhibits superposition. This is the first time such a disentanglement has been achieved in a realistic LLM task to the best of our knowledge..
@@
  • We use these feature dictionaries to design and contextualize evaluations of unsupervised feature dictionaries along the same three axes. Importantly, we aim to develop evaluations that are agnostic to whether or not unsupervised dictionaries use the same concepts as those in the supervised ones.
@@
  • We apply this methodology to feature dictionaries learned by SAEs on either the IOI dataset (task-specific SAEs) or the LLM’s pre-training dataset (full-distribution SAEs). We find that both kinds of SAEs find interpretable features for the task, but task-specific SAEs are able to edit concepts by changing fewer features than full-distribution SAEs. Still, both kinds of SAEs fall short of supervised dictionaries.
@@
  • Finally, we dive deeper into some qualitative phenomena observed in task-specific SAEs: feature occlusion (whereby SAEs have a tendency to learn only the higher-magnitude one of two task-relevant concepts, even when the concepts have slightly different magnitudes), and feature over-splitting (whereby SAEs have a tendency to learn $\gg 2$ features without a clear interpretation for binary concepts). We reproduce both phenomena in simple toy models, suggesting they may generalize beyond our specific setting.
@@
Our results underscore the need for more principled training and evaluation methods in this active area, and suggest that supervised feature dictionaries can be a valuable tool for automating aspects of this process.
@@ 3.1 Overview and Motivation
Step 2: Compute high-quality supervised dictionaries using the attributes. The second step is to learn and validate supervised dictionaries computed using the attribute labels $a_{i}(p),p\sim\mathcal{D}$ (methodology described in Section 4). In particular, not all attribute configurations will result in dictionaries with good approximation, control and interpretability for the task111 To get high-quality supervised dictionaries, we should choose attributes that completely capture the task-relevant information in the prompt, and are moreover compatible with the intermediate states of the model’s computation on the task. Validating that our attributes satisfy this is a non-trivial but key prerequisite for our evaluation to be meaningful. We do this for the IOI task in Section 4, where we show the supervised dictionaries score highly on all tests described in the current section. .
@@ 3.2 Test 1: Sufficiency and Necessity of Dictionary Reconstructions for the Task
Our first test checks if activation reconstructions $\widehat{\mathbf{a}}$ as a whole are sufficient and necessary for the model to perform the task [^2] This test is analogous to the faithfulness/completeness test from Wang et al. (2023); the sufficiency test is also widely used in the literature to evaluate feature dictionaries.. To evaluate sufficiency, we intervene by replacing internal activations $\mathbf{a}$ with their reconstructions $\widehat{\mathbf{a}}$, and measure the drop in performance on the task. To evaluate necessity, we intervene by replacing activations $\mathbf{a}$ with $\mathbb{E}_{\mathcal{D}}\left[\mathbf{a}\right]+\left(\mathbf{a}-\widehat{\mathbf{a}}\right)$, and compare the resulting drop in performance to the effect of replacing activations with $\mathbb{E}_{\mathcal{D}}\left[\mathbf{a}\right]$ alone (also known as mean ablation; see Appendix A.2 for motivation) 333 This test has very few assumptions: it is independent of the attributes we have chosen to describe inputs with, and of any interpretation assigned to the features. However, it only measures the quality of the dictionary as a whole, and not how effectively and sparsely it disentangles any hypothetical concepts in the activations. We include this test because poor performance indicates a fundamental failure of the feature dictionary relative to the task. .
@@ 3.3 Test 2: Sparse Controllability of Attributes
Implementation design space. To achieve independence of interpretations, we frame the problem as a combinatorial optimization problem over the feature dictionary: given an activation $\mathbf{a}_{s}$ for prompt $p_{s}$ and a counterfactual activation $\mathbf{a}_{t}$ for a prompt $p_{t}$ which differs from $p_{s}$ only in the values of the attributes we wish to edit, we can optimize over subsets of features active in $\widehat{\mathbf{a}_{s}}$ to subtract from $\mathbf{a}_{s}$, and subsets of features active in $\widehat{\mathbf{a}_{t}}$ to add to $\mathbf{a}_{s}$. There are multiple optimization objectives possible, such as making the edited activation similar to $\mathbf{a}_{t}$ geometrically, or making model behavior on the edited activation similar to model behavior on $\mathbf{a}_{t}$. The optimization may be constrained by the number of features changed and/or the magnitude of edits 555 There are many possible ways to instantiate such an optimization; in Section 5 we present one such way based on a greedy algorithm minimizing $\ell_{2}$ distance in activation space. Further methodological notes, such as how we implement edits in multiple model components at once, are given in Appendix A.2. . Importantly, this optimization proceeds on a per-prompt basis, which means that it can select different features to edit depending on the prompt being edited. This means it will not disadvantage dictionaries that do not dedicate a uniform w.r.t. all prompts set of features to each attribute (which would be a major assumption to impose).
@@ 4.1 Computing Supervised Feature Dictionaries
Choosing attributes for the IOI task. Not every set of attributes will result in a good approximation of the model’s internal activations; in fact, we find that the choice of attributes is a crucial modeling decision. Recall that, according to Wang et al. (2023), each IOI prompt $p$ is described by three properties influencing how $p$ is processed in the IOI circuit: the subject (S) and indirect object (IO) names, and their relative position (Pos). Motivated by this, we chose the attributes S, IO, and Pos to describe each prompt. We experimented with other choices of attributes, but did not find them to be more successful in our tests (see Appendix A.8 for details888 In particular, we found that another set of attributes, though more expressive in principle, learns to approximate the features we get using the S, IO and Pos attributes through a change-of-variables-like transformation.). We emphasize that there are many other imaginable choices of attributes; see Appendix A.10 for further discussion.
@@ 5.1 Methodology
Figure 4: Trade-offs between edit magnitude and edit success for attribute editing using task-specific SAEs for select cross-sections of the IOI circuit. The x-axis measures the weight removed by an edit (averaged over the attention heads in the cross-section), affine-transformed so that a value of 0 indicates the weight removed by the corresponding supervised edit, and a value of 1 indicates removing $\approx$ all of the mass of the features. The y-axis is an affine transform of the fraction of examples for which the edit results in the same next-token prediction as the ground-truth edit, with a value of 00 corresponding to no intervention, and a value of $1$ corresponding to the supervised edits. The thick lines use our interpretation-agnostic editing method, while the dashed lines choose which features to edit based on $F_{1}$ scores for the corresponding attribute. In both cases, we edit 2, 4 or 6 features (a higher magnitude score indicates editing more features).
@@ Acknowledgements
We thank Sonia Joseph and Can Rager for feedback on earlier versions of this manuscript. We thank the SERI MATS program for making this research possible. A.M. and G.L. were supported by grants from AI Safety Support and Effective Ventures.

### v3

@@ Abstract
Disentangling model activations into meaningful features is a central problem in interpretability. However, the absence of ground-truth for these features in realistic scenarios makes validating recent approaches, such as sparse dictionary learning, elusive. To address this challenge, we propose a framework for evaluating feature dictionaries in the context of specific tasks, by comparing them against supervised feature dictionaries. First, we demonstrate that supervised dictionaries achieve excellent approximation, control, and interpretability of model computations on the task. Second, we use the supervised dictionaries to develop and contextualize evaluations of unsupervised dictionaries along the same three axes.
We apply this framework to the indirect object identification (IOI) task using GPT-2 Small, with sparse autoencoders (SAEs) trained on either the IOI or OpenWebText datasets. We find that these SAEs capture interpretable features for the IOI task, but they are less successful than supervised features in controlling the model. Finally, we observe two qualitative phenomena in SAE training: feature occlusion (where a causally relevant concept is robustly overshadowed by even slightly higher-magnitude ones in the learned features), and feature over-splitting (where binary features split into many smaller, less interpretable features). We hope that our framework will provide a useful step towards more objective and grounded evaluations of sparse dictionary learning methods.
@@ 1 Introduction
While large language models (LLMs) have demonstrated impressive (Vaswani et al., 2017; Devlin et al., 2019; Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) results, the mechanisms behind their successes and failures largely remain a mystery (Olah, 2023). A central problem in this area is how to disentangle internal model representations into meaningful concepts or features. If successful at scale, this research could provide significant scientific and practical value, enabling enhanced model robustness, controllability, interpretability, and debugging (Gandelsman et al., 2023; Nanda et al., 2023; Marks et al., 2024).
@@
A leading hypothesis for how LLMs represent and use features is the linear representation hypothesis (Mikolov et al., 2013b; Grand et al., 2018; Li et al., 2021; Abdou et al., 2021; Nanda et al., 2023). A strong version of this hypothesis posits that individual activations of a model can be decomposed into sparse linear combinations of features from a large, shared feature dictionary. Recently, a series of works has proposed applying the (unsupervised) sparse autoencoder (SAE) framework to find such dictionaries (Olshausen & Field, 1997; Faruqui et al., 2015; Goh, 2016; Arora et al., 2018; Yun et al., 2021; Cunningham et al., 2023; Bricken et al., 2023).
Figure 1: Overview of our evaluation pipeline. We begin by selecting a specific model capability and then disentangling model activations into capability-relevant features using supervision. Then, we evaluate a given feature dictionary w.r.t. this capability, using the supervised features as a benchmark. We test the extent to which (1) the feature dictionary’s reconstructions of the activations are necessary and sufficient for the capability, (2) the features can be used to edit capability-relevant information in internal model representations (agnostic of feature interpretations), and (3) the features can be interpreted w.r.t. the capability in a manner consistent with their causal role.
@@
While there are promising initial results (Bricken et al., 2023), this research area faces a key obstacle: we cannot directly evaluate the usefulness of features learned by an SAE, as we do not know the hypothetical ‘true’ features to begin with; indeed, finding them is the reason we use SAEs in the first place. This has two consequences: (1) the training objective of an SAE – balancing $\ell_{2}$ reconstruction with an $\ell_{1}$ penalty on feature coefficients – is only a substitute for the true goal of finding meaningful features, and (2) the metrics used to evaluate SAEs are indirect, relying on proxies for the features, success in toy models, or non-trivial assumptions on SAE learning (Elhage et al., 2022b; Bricken et al., 2023; Sharkey et al., 2023). This hampers progress in the field, as there is no simple, objective and direct way to compare trained SAEs. In this paper, we take steps towards addressing this challenge. Our contributions are outlined as follows:
@@
  • We propose a principled method for computing sparse feature dictionaries to decompose language model activations on realistic tasks, using supervision from attributes describing the prompts.
@@
  • We apply this method to the IOI task, demonstrating that these dictionaries exhibit three desirable properties in the task’s context: (1) sufficiency and necessity of activation reconstructions, (2) sparse controllability of model behavior via feature editing, and (3) interpretability of the features consistent with their causal role111 As a by-product, our supervised feature dictionaries also demonstrate that task activations can be usefully disentangled in a way that adheres to the linear representation hypothesis and exhibits superposition. To the best of our knowledge, this is the first time such a disentanglement has been achieved in a realistic LLM task..
@@
  • We use these feature dictionaries to design and contextualize evaluations of unsupervised feature dictionaries along the same three axes. Importantly, we aim to develop evaluations that are agnostic to whether the unsupervised dictionaries use the same concepts as the supervised ones.
@@
  • We apply this methodology to feature dictionaries learned by SAEs trained on either the IOI dataset (task-specific SAEs) or the LLM’s pre-training dataset (full-distribution SAEs). We find that both types of SAEs find interpretable features for the task, but task-specific SAEs allow us to edit attributes by changing fewer features compared to full-distribution SAEs. However, both kinds of SAEs fall short of supervised dictionaries.
@@
  • Finally, we dive deeper into some qualitative phenomena observed in task-specific SAEs: feature occlusion (where SAEs have a tendency to learn only the higher-magnitude of two task-relevant attributes, even when the attributes have slightly different magnitudes), and feature over-splitting (where SAEs tend to learn $\gg 2$ features without a clear interpretation for binary concepts). We reproduce both phenomena in simple toy models, suggesting they may generalize beyond our specific setting.
@@
Our results underscore the need for more principled training and evaluation methods in this active area and suggest that supervised feature dictionaries can be a valuable tool for automating aspects of this process.
@@ 3.1 Overview and Motivation
Step 2: Compute high-quality supervised dictionaries using the attributes. The second step is to learn and validate supervised dictionaries computed using the attribute labels $a_{i}(p),p\sim\mathcal{D}$ (methodology described in Section 4). In particular, not all attribute configurations will result in dictionaries with good approximation, control and interpretability for the task222 To get high-quality supervised dictionaries, we should choose attributes that completely capture the task-relevant information in the prompt, and are moreover compatible with the intermediate states of the model’s computation on the task. Validating that our attributes satisfy this is a non-trivial but key prerequisite for our evaluation to be meaningful. We do this for the IOI task in Section 4, where we show the supervised dictionaries score highly on all tests described in the current section. .
@@ 3.2 Test 1: Sufficiency and Necessity of Dictionary Reconstructions for the Task
Our first test checks if activation reconstructions $\widehat{\mathbf{a}}$ as a whole are sufficient and necessary for the model to perform the task [^3] This test is analogous to the faithfulness/completeness test from Wang et al. (2023); the sufficiency test is also widely used in the literature to evaluate feature dictionaries.. To evaluate sufficiency, we intervene by replacing internal activations $\mathbf{a}$ with their reconstructions $\widehat{\mathbf{a}}$, and measure the drop in performance on the task. To evaluate necessity, we intervene by replacing activations $\mathbf{a}$ with $\mathbb{E}_{\mathcal{D}}\left[\mathbf{a}\right]+\left(\mathbf{a}-\widehat{\mathbf{a}}\right)$, and compare the resulting drop in performance to the effect of replacing activations with $\mathbb{E}_{\mathcal{D}}\left[\mathbf{a}\right]$ alone (also known as mean ablation; see Appendix A.2 for motivation) 444 This test has very few assumptions: it is independent of the attributes we have chosen to describe inputs with, and of any interpretation assigned to the features. However, it only measures the quality of the dictionary as a whole, and not how effectively and sparsely it disentangles any hypothetical concepts in the activations. We include this test because poor performance indicates a fundamental failure of the feature dictionary relative to the task. .
@@ 3.3 Test 2: Sparse Controllability of Attributes
Implementation design space. To achieve independence of interpretations, we frame the problem as a combinatorial optimization problem over the feature dictionary: given an activation $\mathbf{a}_{s}$ for prompt $p_{s}$ and a counterfactual activation $\mathbf{a}_{t}$ for a prompt $p_{t}$ which differs from $p_{s}$ only in the values of the attributes we wish to edit, we can optimize over subsets of features active in $\widehat{\mathbf{a}_{s}}$ to subtract from $\mathbf{a}_{s}$, and subsets of features active in $\widehat{\mathbf{a}_{t}}$ to add to $\mathbf{a}_{s}$. There are multiple optimization objectives possible, such as making the edited activation similar to $\mathbf{a}_{t}$ geometrically, or making model behavior on the edited activation similar to model behavior on $\mathbf{a}_{t}$. The optimization may be constrained by the number of features changed and/or the magnitude of edits 666 There are many possible ways to instantiate such an optimization; in Section 5 we present one such way based on a greedy algorithm minimizing $\ell_{2}$ distance in activation space. Further methodological notes, such as how we implement edits in multiple model components at once, are given in Appendix A.2. . Importantly, this optimization proceeds on a per-prompt basis, which means that it can select different features to edit depending on the prompt being edited. This means it will not disadvantage dictionaries that do not dedicate a uniform w.r.t. all prompts set of features to each attribute (which would be a major assumption to impose).
@@ 4.1 Computing Supervised Feature Dictionaries
Choosing attributes for the IOI task. Not every set of attributes will result in a good approximation of the model’s internal activations; in fact, we find that the choice of attributes is a crucial modeling decision. Recall that, according to Wang et al. (2023), each IOI prompt $p$ is described by three properties influencing how $p$ is processed in the IOI circuit: the subject (S) and indirect object (IO) names, and their relative position (Pos). Motivated by this, we chose the attributes S, IO, and Pos to describe each prompt. We experimented with other choices of attributes, but did not find them to be more successful in our tests (see Appendix A.8 for details999 In particular, we found that another set of attributes, though more expressive in principle, learns to approximate the features we get using the S, IO and Pos attributes through a change-of-variables-like transformation.). We emphasize that there are many other imaginable choices of attributes; see Appendix A.10 for further discussion.
@@ 5.1 Methodology
Figure 4: Trade-offs between edit magnitude and edit success for attribute editing using task-specific SAEs for select IOI circuit cross-sections. The x-axis measures the weight (see Subsection 5.1) of the features removed by the edit (features added are not reported in this plot), averaged over the attention heads in the cross-section. This metric is affine-transformed so that a value of 0 indicates the weight removed by the corresponding supervised edit, and a value of 1 indicates that the edit removed all features in the reconstruction. The y-axis is an affine transform of the fraction of examples for which the edit results in the same next-token prediction as the ground-truth edit, with a value of 00 corresponding to no intervention, and a value of $1$ corresponding to the supervised edits. Results for our interpretation-agnostic/interpretation-aware editing methods are shown as thick/dashed lines respectively. For both methods, we edit 2, 4 or 6 features (a higher magnitude score indicates editing more features).
@@ Acknowledgements
We thank Sonia Joseph and Can Rager for feedback on earlier versions of this manuscript. We thank the SERI MATS program for making this collaboration possible. We also thank Atticus Geiger for valuable discussions and supervision during some early versions of this project. AM was supported by grants from AI Safety Support and Effective Ventures. GL was supported by grants from AI Safety Support.